# CrochetBench: Can Vision-Language Models Move from Describing to Doing in Crochet Domain?

### v1

@@ Abstract
We present CrochetBench, a benchmark for evaluating the ability of multimodal large language models to perform fine-grained, low-level procedural reasoning in the domain of crochet. Unlike prior benchmarks that focus on high-level description or visual question answering, CrochetBench shifts the emphasis from describing to doing: models are required to recognize stitches, select structurally appropriate instructions, and generate compilable crochet procedures. We adopt the CrochetPARADE DSL as our intermediate representation, enabling structural validation and functional evaluation via execution. The benchmark covers tasks including stitch classification, instruction grounding, and both natural language and image-to-DSL translation. Across all tasks, performance sharply declines as the evaluation shifts from surface-level similarity to executable correctness, exposing limitations in long-range symbolic reasoning and 3D-aware procedural synthesis. CrochetBench offers a new lens for assessing procedural competence in multimodal models and highlights the gap between surface-level understanding and executable precision in real-world creative domains. Code is available at https://github.com/Peiyu-Georgia-Li/crochetBench.
@@ 1 Introduction
Procedural crafts such as crochet present a distinctive frontier for multimodal learning. Unlike traditional captioning or recipe datasets (chefFusion; Hu_2023_ICCV; mohbat2024llava), crochet patterns intertwine three interdependent modalities: (i) structured symbolic language, where stitch abbreviations and counts define a precise grammar of construction; (ii) long-form natural language, which provides contextual guidance such as materials and sizing; and (iii) visual evidence, including photographs of completed objects and motif diagrams. Success requires not just alignment across modalities but step-wise reasoning that preserves procedural fidelity, making the challenge closer to program synthesis than generic description.
Crochet also offers a unique testbed for 3D-aware reasoning. Each stitch encodes both local geometry and global connectivity, forming a topological structure that must be preserved across steps. Generating or interpreting patterns thus demands reasoning over how sequential operations accumulate into volumetric form. In effect, crochet couples symbolic instruction following with embodied spatial reasoning, cultivating abilities essential for domains where language must ground into physical tasks.
Despite the rapid growth of multimodal benchmarks (fu2024mmecomprehensiveevaluationbenchmark; li2023seed; zhang-etal-2025-ngqa; yue2024mmmu), existing datasets have largely focused on description or grounding. COCO (lin2015microsoftcococommonobjects) catalyzed captioning research, TextCaps (sidorov2020textcapsdatasetimagecaptioning) extended it to text-in-the-wild, and Recipe1M (marin2019recipe1mdatasetlearningcrossmodal) explored cross-modal cooking instructions. While recipes also involve multi-step procedures, validating correctness typically requires real-world execution, making large-scale evaluation slow and resource-intensive. Crochet, by contrast, provides a symbolic domain where outputs can be automatically verified through DSL compilation, enabling scalable and efficient study of step-wise reasoning. For a more detailed survey, see Appendix A. Yet these benchmarks stop short of testing whether models can follow symbolic grammars, respect numerical and spatial constraints, and produce outputs that are executable. Current systems can describe, but not reliably do.
CrochetBench fills this gap by centering evaluation on instructional fidelity: can models not only recognize and generate, but also output step-wise, compilable instructions that respect symbolic, numerical, and topological structure? Each example in CrochetBench is a multimodal package—structured JSON metadata (stitch inventories and abbreviations), full-text procedures with rows/rounds and conditionals, and paired images of finished objects and motifs. Crucially, CrochetBench is paired with CrochetPARADE (crochetparade), a domain-specific language (DSL) enabling executable evaluation, where natural language instructions are translated into compilable code enforcing geometric and topological coherence.
Our contributions are fourfold: (1) CrochetBench, the first executable benchmark for procedural crafts, unifying symbolic, textual, and visual modalities with evaluation protocols emphasizing procedural fidelity and 3D-aware reasoning; (2) a comprehensive task suite spanning recognition, comprehension, generation, and DSL translation; (3) integration of CrochetPARADE into an executable pipeline, enabling scalable, automated verification of outputs—unlike domains such as cooking, which require real-world execution—thereby shifting evaluation from surface similarity to procedural fidelity; and (4) baseline analyses of state-of-the-art VLMs/MLLMs, revealing systematic weaknesses including hallucinations, captioning bias, and structural artifacts.
Taken together, CrochetBench opens a new direction for multimodal research: moving beyond describing what we see, toward generating executable procedures that respect symbolic grammar, numerical accuracy, and topological coherence—paving the way for models that can reason in structured 3D spaces.
@@ 2 Dataset Description
CrochetBench is a large-scale, structured benchmark consisting of 6,085 crochet patterns spanning 55 distinct project categories. It was constructed by collecting publicly available patterns from the Yarnspirations website [^1] https://www.yarnspirations.com/collections/patterns, a widely used repository for fiber arts. The raw patterns were originally formatted as PDF documents, which were parsed and normalized using a GPT-4o-mini-based conversion pipeline. This process extracted and standardized key fields including pattern metadata, materials, measurements, gauge, abbreviations, and full step-by-step instructions. Each entry was converted into a machine-readable JSON object with a consistent schema, and 98.77% of patterns include an associated product image, enabling multimodal supervision.
Figure 2: Skill level distribution across the CrochetBench dataset. Note that the “Experienced” slice (1.3%) is annotated externally due to its small size.
The dataset supports diverse real-world crochet practices, with project types ranging from simple accessories to complex garments. Figure 1 lists the ten most common categories by frequency. The majority of patterns belong to a small number of dominant types—Afghans and Blankets alone account for over one-quarter of the dataset. More details can be found at Appendix B.
Each pattern is labeled with one of four primary skill levels—beginner, easy, intermediate, or experienced. This allows for stratified evaluation across complexity tiers. Figure 2 shows the skill level distribution, which is strongly skewed toward beginner-friendly content. Only one pattern (0.02%) is missing a skill level label. More details can be found at Appendix B.
Instructional complexity varies substantially across patterns. The number of characters in each instruction ranges from 20 to over 30,000, with a mean of 3,216 and a median of 2,453. Abbreviation counts (i.e., unique stitch tokens per pattern) range from 1 to 31, with an average of 10.6. These statistics are summarized in Appendix B.1. We observe a clear correlation between skill level and instruction length: beginner patterns tend to be short and use fewer abbreviations, while experienced patterns are significantly longer and more symbolically dense.
In addition to symbolic complexity, the dataset contains 3,143 abbreviation instances mapped to 789 unique standardized stitch tokens. This lexical mapping enables tasks such as vocabulary translation, sequence generation, and instruction validation. Beyond raw instructions, the structured schema also records rich metadata, including gauge, hook size, yarn weight, and measurements. A representative dataset entry is provided in Appendix 10.
Overall, CrochetBench provides a rich resource for multimodal modeling, symbolic reasoning, and structure-aware generation. Its coverage across diverse categories and complexity levels enables broad benchmarking of both open-ended generation and instruction fidelity tasks.
Table 1: Overall statistics of the CrochetBench dataset.
Total Patterns Image Coverage Avg. Instr. Length #Project Types
CrochetBench 6,085 98.77% 3,216 characters 55
3 Experiment
We empirically evaluate CrochetBench by defining a set of structured tasks and benchmarking a diverse pool of multimodal large language models. Our experiments are designed to test models across progressively challenging stages, from low-level recognition to high-level executable synthesis.
Benchmark Tasks.
Table 2 summarizes the four evaluation tasks in CrochetBench, which progress systematically from recognition to comprehension, generation, and ultimately executable synthesis. Task A (Stitch Recognition) evaluates a model’s ability to detect symbolic primitives in crochet images, establishing the foundation for multimodal perception. Task B (Instruction Selection) requires models to align visual evidence with candidate textual instructions, thereby testing multimodal grounding and fine-grained comprehension. Unlike conventional description tasks, the candidates are procedural steps rather than captions, and correct selection often requires reasoning about how local steps contribute to the final product. Task C (Instruction Generation) advances from comprehension to open-ended production, challenging models to generate natural-language procedural instructions that are both perceptually grounded and linguistically faithful to domain conventions. Evaluation here emphasizes lexical and symbolic fidelity, but does not directly capture structural validity. Task D (Instruction-to-DSL Translation) addresses this gap by requiring models to output a compilable program in a domain-specific language (DSL). The step-level variant tests local semantic grounding, while the project-level variant demands global structural consistency across the entire pattern. Compilation-based evaluation directly measures executable faithfulness, ensuring that generated instructions are not only linguistically plausible but also structurally sound. Together, these tasks form a structured evaluation ladder that moves from perceptual recognition to programmatic execution, probing both low-level perception and high-level symbolic reasoning.
Table 2: Overview of benchmark tasks in CrochetBench. Tasks progress from recognition to comprehension, generation, and executable synthesis.
@@ 3 Experiment
ID Ability Tested Task Evaluation Metrics Test Size
A Recognition Stitch Recognition F1, Precision, Recall
6,009
(CrochetBench-A)
B Comprehension Instruction Selection Accuracy
6,003
(CrochetBench-B)
C Generation Instruction Generation BLEU, ROUGE, ChrF
6,009
(CrochetBench-C)
@@ C Generation Instruction Generation BLEU, ROUGE, ChrF
D Formalization Instr.-to-DSL (Step) Valid Pattern Rate
119
(CrochetBench-Dstep)
@@
Instr.-to-DSL (Project) Valid Pattern Rate
100
(CrochetBench-Dproj)
Model Selection.
We evaluate a diverse set of vision–language models (VLMs) spanning both open-source and closed-source families. The open-source group includes Salesforce BLIP-2 Flan-T5 XL (3B), a perception-focused baseline widely used in image–text tasks; Google Gemma 3 (4B) and Qwen2-VL (7B), two recent models trained with large-scale multimodal alignment; and DeepSeek-VL (7B), a larger open-source model designed for stronger vision–language reasoning. On the closed-source side, we assess GPT-4o, Gemini 2.5 Flash-Lite, and Claude Sonnet 4, which represent state-of-the-art commercial VLMs that push the limits of multimodal reasoning. This selection spans perception-heavy baselines to general-purpose multimodal reasoning systems, enabling us to evaluate capabilities across both research-grade and production-grade settings. [^2] We use the term “VLM” broadly to include both traditional vision–language models (e.g., BLIP-2) and modern multimodal large language models (e.g., GPT-4o, Gemini, Claude).
@@ 3.1 Task A: Stitch Recognition
The first task evaluates a model’s ability to identify crochet stitch types from an image of a finished product. We construct CrochetBench-A, a subset of 6,009 examples from the full benchmark, where each product image is paired with ground-truth stitch annotations. These labels are derived from the official pattern instructions and normalized into a standardized set of stitch abbreviations (e.g., sc, hdc, dc) to ensure consistency across patterns. Unlike standard image classification, this is a multi-label prediction problem: multiple stitches may co-occur within the same image, often with subtle visual differences in texture and geometry. This task therefore probes fine-grained visual grounding of structured crochet semantics.
@@
For each example, we compute overlap between the predicted and reference stitch sets. True Positives (TP) are stitches correctly predicted; False Positives (FP) are stitches predicted but not in the reference; and False Negatives (FN) are stitches in the reference but missed by the model. From these counts, we compute precision (fraction of correct predictions among all predictions), recall (fraction of ground-truth stitches recovered), and F1 score (harmonic mean). Metrics are averaged across examples to provide overall performance. This formulation rewards models that recover all present stitches while avoiding spurious predictions.
@@
Table 3: Evaluation results on the Stitch Recognition task. We report Precision, Recall, and F1. Best results are bold; second-best are underlined.
Model Size Precision Recall F1
Open Source Salesforce BLIP-2 Flan-T5 XL 3B 0.2953 0.2303 0.2250
Google Gemma 3 4B 0.2054 0.1021 0.1265
DeepSeek-VL 7B 0.5447 0.7476 0.6060
Qwen2-VL 7B 0.5414 0.6974 0.5816
Closed Source GPT-4o 0.6214 0.5939 0.5801
Gemini 2.5 Flash-Lite 0.7449 0.4977 0.5683
Claude Sonnet 4 0.7861 0.5312 0.6094
Results. Claude Sonnet 4 achieves the best overall F1 score, demonstrating strong precision in stitch recognition. Among open-source models, DeepSeek-VL performs best, with notably high recall, while Qwen2-VL offers a competitive balance. These results highlight the gap between commercial VLMs and open-source alternatives, but also suggest that large open-source models are beginning to approach closed-source performance in fine-grained recognition.
@@ 3.2 Task B: Instruction Selection
The second task evaluates whether a model can correctly align a finished crochet product image with its corresponding natural-language instruction. We construct CrochetBench-B, a subset of 6,003 examples, where each item consists of one correct instruction and three distractor instructions sampled from the same project category (e.g., hats, rugs). Distractors are carefully chosen to share structural motifs and vocabulary, thereby increasing difficulty and requiring fine-grained visual–textual alignment rather than reliance on superficial cues. The answer distribution across options is approximately uniform (A: 24.9%, B: 25.7%, C: 23.7%, D: 25.7%).
@@
To enable scalable assessment, we formulate the task as a 4-way multiple-choice question (MCQ). The model must select one option (A–D), with exactly one correct answer. Predictions are extracted using a lightweight regex-based method that identifies explicit letter outputs (e.g., “A”, “The answer is B”). If no parsable choice is found, the response is marked as unanswered. Accuracy is reported as the evaluation metric.
This task provides a controlled measure of visual grounding and semantic alignment between image content and procedural text, without requiring generative modeling. It highlights the challenge of distinguishing between subtle visual cues and domain-specific terminology, which is essential for bridging perception and structured instruction understanding.
Table 4: Evaluation results on the Instruction Selection task (4-way multiple choice). We report Accuracy. Best results are bold; second-best are underlined.
Model Size Accuracy
Open Source Salesforce BLIP-2 Flan-T5 XL 3B 0.2562
Google Gemma 3 4B 0.2494
DeepSeek-VL 7B 0.2892
Qwen2-VL 7B 0.4196
Closed Source GPT-4o 0.5811
Gemini 2.5 Flash-Lite 0.5563
Claude Sonnet 4 0.5739
Results. GPT-4o achieves the highest accuracy, closely followed by Claude and Gemini, all of which substantially outperform open-source VLMs. Among open-source models, Qwen2-VL is the strongest, while BLIP-2 and Gemma remain near chance level, underscoring the difficulty of visually grounded instruction matching in this domain.

### v2

@@ Abstract
While multimodal large language models can describe visual content, their ability to generate executable procedures remains underexplored. CrochetBench presented in this paper evaluates this shift from describing to doing through fine-grained procedural reasoning in crochet: models must recognize stitches, select structurally appropriate instructions, and generate compilable procedures. We adopt the CrochetPARADE DSL as our intermediate representation, enabling structural validation and functional evaluation via execution. The benchmark covers tasks including stitch classification, instruction grounding, and both natural language and image-to-DSL translation. Across all tasks, performance sharply decreases as the evaluation shifts from surface-level similarity to executable correctness, revealing limitations in long-range symbolic reasoning and 3D-aware procedural synthesis. Our proposed CrochetBench offers a new lens for assessing procedural competence in multimodal models and highlights the gap between surface-level understanding and executable precision in real-world creative domains. Code is available at https://anonymous.4open.science/r/crochet-82E6/README.md.
CrochetBench: Can Vision-Language Models Move from Describing to Doing in Crochet Domain?
Peiyu Li*, Xiaobao Huang*, Ting Hua, Nitesh V. Chawla University of Notre Dame Notre Dame, IN 46556 USA {pli9, xhuang2, thua, nchawla}@nd.edu
**footnotetext: These authors contributed equally.
Figure 1: End-to-end data construction and benchmarking workflow for CrochetBench. The left panel illustrates the data collection pipeline: we download PDF files and image links from crochet pattern websites, and parse them using GPT-4o-mini to produce structured JSON files. From each JSON record, we derive four supervised datasets: (A) stitch-level labels, (B) multiple-choice instruction selection items, (C) natural-language instruction generation pairs, and (D) expert-annotated DSL programs for procedural synthesis. The right panel summarizes the four benchmarking tasks: Task A evaluates multi-label visual stitch recognition; Task B measures vision-to-instruction grounding via MCQ selection; Task C assesses vision-conditioned natural-language instruction generation; and Task D tests stateful procedural reasoning via NL-to-DSL (Natural Language to Domain-Specific Language) translation with execution-based metrics.
@@ 1 Introduction
Multimodal large language models can describe what they see, but can they explain how to make it? Given an image of a piece of furniture or a crocheted blanket, current models can fluently describe its appearance, yet cannot generate the step-by-step instructions needed to assemble or recreate it. Closing this gap, from perception to executable procedure, is critical for applications in robotics, CAD design, and procedural content generation. However, evaluating procedural competence is difficult: most real-world procedures require physical execution to verify correctness. One domain offers a rare exception: crochet instruction synthesis. These crochet instructions follow a formal symbolic grammar that can be compiled and validated automatically, enabling scalable evaluation of procedural competence without requiring costly physical execution.
Existing multimodal benchmarks largely focus on descriptive tasks. Image captioning benchmarks such as COCO (Lin et al., 2014) and TextCaps (Sidorov et al., 2020) evaluate whether models can generate natural language descriptions of visual scenes. Text-to-code benchmarks such as HumanEval (Chen, 2021a) evaluate program synthesis but lack visual grounding, while image-to-code benchmarks such as pix2code (Beltramelli, 2017) test layout generation rather than stateful procedural reasoning. Even procedural datasets such as Recipe1M+ (Marin et al., 2018) primarily assess cross-modal alignment rather than executable correctness, since validating a recipe requires cooking it, a slow and resource-intensive process. No existing benchmark combines visual understanding, multi-step procedural decomposition, and automated executable validation.
To address this gap, we turn to crochet: an ideal testbed for visual procedural reasoning. Each crochet pattern presents a vision-to-procedure challenge. Given an image of a finished artifact, models must infer the sequential construction process: which stitches were used, in what order, and with what counts. This requires (1) fine-grained visual recognition of stitch types from texture and topology, (2) procedural decomposition into ordered operations with symbolic grounding, (3) stateful reasoning to track counts across steps, and (4) constraint satisfaction to ensure topological consistency. What makes crochet unique is that, crochet instructions can be formalized into domain-specific languages and validated through compilation, which enables the automated executable evaluation that other procedural domains lack.
In this paper, we introduce CrochetBench, the first benchmark for executable visual procedural reasoning in creative domains. Our CrochetBench comprises 6,085 crochet patterns spanning 55 categories with over 98% image coverage, ranging from beginner to expert complexity. We design four tasks of increasing difficulty: Task A (Stitch Recognition) evaluates multi-label visual classification; Task B (Instruction Selection) evaluates vision-to-language grounding via multiple-choice matching; Task C (Instruction Generation) evaluates free-form instruction synthesis from images; and Task D (DSL Translation) evaluates translation into executable CrochetPARADE programs (Tassev, 2025). Together, these tasks shift evaluation from surface-level fluency to procedural correctness, whether outputs compile and preserve structural constraints.
We evaluate nine state-of-the-art vision-language models on CrochetBench, including GPT-4o (Hurst et al., 2024), Gemini 2.5 Flash-Lite (Comanici et al., 2025), Claude Sonnet 4 (Anthropic, 2025), and open-source alternatives spanning 3B to 72B parameters.
Our evaluation reveals that current models exhibit a consistent gap between perception and procedural synthesis. Models achieve moderate performance on stitch recognition and instruction selection, but struggle to generate multi-step instructions or executable DSL programs. A case study (Figure 2) illustrates this pattern: even when models produce fluent instructions that correctly identify local details such as colors and motifs, the resulting procedures misconstruct global geometry and fail under execution. Also, neither model scaling nor supervised finetuning closes this gap. Larger models often achieve lower valid pattern rates during execution due to a tendency to invent novel stitch types not defined in the DSL, while finetuning improves text-based metrics without improving executable correctness.
Our contributions are threefold: (1) We release CrochetBench, the first executable benchmark for procedural textile crafts, comprising 6,085 patterns integrated with the CrochetPARADE DSL for automated validation beyond text similarity. (2) We design four tasks of increasing difficulty—stitch recognition, instruction selection, instruction generation, and DSL translation—that enable fine-grained diagnosis of procedural competence. (3) We evaluate nine state-of-the-art vision–language models on CrochetBench and reveal a consistent gap between recognition and executable synthesis that persists across model scales and finetuning.
@@ 2 Dataset Description
CrochetBench is a large-scale, structured benchmark comprising 6,085 crochet patterns across 55 distinct project categories. As shown in the left panel of Figure 1, the dataset is constructed from publicly available patterns on the Yarnspirations website***https://www.yarnspirations.com/collections/patterns, a widely used repository in the fiber-arts community. The raw patterns (originally distributed as PDF documents) were parsed and normalized through a GPT-4o-mini–based conversion pipeline that extracted and standardized key fields such as metadata, materials, measurements, gauge, abbreviations, and full step-by-step instructions. Each pattern was then transformed into a JSON object following a consistent schema. Each pattern pairs a finished-product image with full natural-language instructions, enabling both perception- and generation-based evaluation.
CrochetBench exhibits substantial diversity in both visual appearance and procedural complexity. Patterns vary widely in instruction length, stitch vocabulary, and structural depth, supporting evaluation across different levels of difficulty and long-horizon reasoning. Detailed statistics, skill-level distributions, and a JSON example are provided in the Appendix B.
3 Tasks
A central goal of CrochetBench is to evaluate whether multimodal LLMs can move beyond surface-level visual description and produce procedurally correct crochet instructions. As shown in Table 1, CrochetBench is organized as a progression of four tasks that isolate the core cognitive abilities required for real-world crochet reasoning.
Tasks A and B focus on perception and comprehension, representing the minimum prerequisites for procedural understanding. Stitch recognition and instruction selection evaluate whether models can identify stitch types from images and correctly associate visual patterns with corresponding natural-language instructions. However, accomplishing these two tasks does not guarantee the ability to synthesize a valid crochet procedure. Tasks C and D therefore target procedural generation and formalization, requiring models to produce coherent, stepwise natural-language instructions or executable CrochetPARADE programs. These tasks demand the integration of visual grounding, temporal consistency, symbolic manipulation, and domain-specific constraints. The following subsections describe each task in detail.
@@ 3 Tasks
ID Ability Tested Task Evaluation Metrics Test Size Eval Set
A Recognition Stitch Recognition F1, Precision, Recall 6,009 CrochetBench-A
B Comprehension Instruction Selection Accuracy 6,003 CrochetBench-B
C Generation Instruction Generation BLEU, ROUGE, ChrF 6,009 CrochetBench-C
@@
D Formalization Instr.-to-DSL (Step) Valid Pattern Rate 119 CrochetBench-Dstep
@@
Instr.-to-DSL (Project) Valid Pattern Rate, Dino Similarity 100 CrochetBench-Dproj
Table 1: CrochetBench task overview and evaluation metrics.
@@ 3.1 Task A: Stitch Recognition
Task A evaluates a model’s ability to identify crochet stitch types from an image of a finished product. We construct CrochetBench-A, a subset of 6,009 examples from the full benchmark, where each product image is paired with ground-truth stitch annotations. These labels are derived from the official pattern instructions and normalized into a standardized set of stitch abbreviations (e.g., sc, hdc, dc) to ensure consistency across patterns. Unlike standard image classification, this is a multi-label prediction problem: multiple stitches may co-occur within the same image, often with subtle visual differences in texture and geometry. This task therefore probes fine-grained visual grounding of structured crochet semantics.
@@
For each example, we compute overlap between the predicted and reference stitch sets. True Positives (TP) are stitches correctly predicted; False Positives (FP) are stitches predicted but not in the reference; and False Negatives (FN) are stitches in the reference but missed by the model. From these counts, we compute precision (fraction of correct predictions among all predictions), recall (fraction of ground-truth stitches recovered), and F1 score (harmonic mean) Powers (2020). Metrics are averaged across examples to provide overall performance. This formulation rewards models that recover all present stitches while avoiding spurious predictions.
@@ 3.2 Task B: Instruction Selection
Task B evaluates whether a model can correctly associate an image of a finished crochet product image with its corresponding natural-language instruction. We construct CrochetBench-B, a subset of 6,003 examples, where each instance contains one ground-truth instruction and three distractor instructions sampled from the same project category (e.g., hats, rugs). Because distractors originate from the same category, they share similar visual and lexical structure, thereby increasing task difficulty and preventing solutions based on superficial lexical overlap. The answer distribution across options is approximately uniform (A: 24.9%, B: 25.7%, C: 23.7%, D: 25.7%), ensuring no positional bias.
@@
To support scalable and reproducible benchmarking, we formulate the task as a four-way multiple-choice question (MCQ). The model must select one option (A–D), with exactly one correct answer. Predictions are extracted using a deterministic regex-based parser that identifies explicit letter-based responses (e.g., “A”, “Option B”, “The answer is D”). Accuracy is used as the evaluation metric.
This task provides a controlled measure of visual grounding and semantic alignment between images and procedural text, without requiring free-form generation. By forcing discrimination among near-neighbor instructions, Task B probes whether models can leverage fine-grained visual cues and domain-specific stitch semantics, which are essential precursors to reliable procedural instruction generation.